Greenhab Report – April 3rd

GreenHab Officer: Erika Rydberg

Environmental control:

Ambient, door open.

Shade cloth (80%) on

Average temperatures:

Average: 25.5 C

Low: 20 C
High: 31.4 C

Hours of supplemental light: 0

Daily water usage for crops: 5 gallons

Water in Blue Tank— 203 gallons

Time(s) of watering for crops: 8:45 and 17:30

Changes to crops: Continued to remove new basil flowers to keep basil growing.

Narrative: It warmed up in the Greenhab today, kept a close eye on the temperature.

Harvest: None

Support/supplies needed: None

Astronomy Report – April 3rd

Astronomy Report

Name: Amanda Manget

Crew: 208
Date: 03-APR-2019

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-WF

Objects to be Imaged this Evening (if weather permits):

Submitted March 27th: Pinwheel (active 7/10 images taken)

Submitted April 1st: Crab Nebula, Dumbbell Nebula

Objects Already Imaged: Pinwheel with MDRS-14, Whirlpool Galaxy with MDRS-WF

Problems Encountered: High winds last night causing dome to stay closed. If weather permits, images will be taken tonight.

MUSK OBSERVATORY

Solar Features Observed: N/A

Problems Encountered: N/A

Journalist Report – April 3rd

[title Journalist Report – April 3rd]

Crew 208 Journalist Report

Date: April 3rd 2019

Author: Erika Rydberg

Making It Work On Mars

Today we had a number of successes in our experience and travels here on Mars. We successfully printed another 3D map, we were able to fix our non-functioning 3D printer, and Dean and myself had solid one-hour rides on our bicycle to charge the battery we’re using to power our 3D printer.

David, Amanda, and I had an awesome EVA out to check out an area north of the Hab, where the landscape is called Grey Moon. Did it look moon-like? Well we haven’t visited there yet ourselves, but it was surprising and the color description was accurate. We hoped to find fossils, but instead found only the breathtaking landscape in front of us, which was enough to satisfy us on our journey. The longer we stay, the more the landscape continues to be surprising, surreal, and varied.

A short report for today, but as usual we’re never quite sure what tomorrow will hold.

Sol Summary – April 3rd

Crew 208 Sol Summary Report 03-APR-2019

Sol: 6

Summary Title: Geology Rocks! (in sim)

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

Crew 208 continued 3D printing in the Hab and worked on their research manuscripts.

Useful things to 3D print at MDRS:

Ice cube tray

ENG + HSO + GHO used the Curiosity and Spirit rovers during their EVA to search for fossils.

Crew XO + HSO tried out our Space VR goggles.

Crew XO + GHO each exercised for 1 hour to generate power for our bike-powered 3D printing study.

Look Ahead Plan: Continue 3D printing in Hab and data collection for thoracentesis study in Science Dome. Plan to break sim on Sol 8.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Nominal

EVA: #5 to Grey Moon and Beige Moon

Reports to be filed: Operations, Greenhab, Journalist, EVA Report, EVA Request

Support Requested: None

Operations Report – April 2nd

Crew 208 Operations Report 02-APR-2019

SOL: 5

Name of person filing report: Amanda Manget

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 14

Solar –SOC 81% @ 1530 (before David turned on generator for the night)

Diesel Reading – 46%

Propane Reading – 52%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – water before transfer: 96.81, water after: 214.37

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 208 gallons

Static to Loft Pump used – Yes

Water Meter: 0142368.7 gallons @ 18:12

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours: 85.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 55.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 85.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where? David delivered water using this car, then Dean and I transferred it into the tank.

CrewCar used and why, where? Not Used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read > 12.5 volts

Summary of Hab operations: Checked SOC of electrical system as follows:

0730 SOC 100%, generator turned off by David

12:20 SOC 89%, generator off

15:30 SOC 81%, generator turned on by David

17:30 SOC 81%, generator on

18:50 SOC 81%, generator on

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: None

Cheers,

Julielynn Wong, MD, MPH, FACPM
Chief Medical Maker

Astronomy Report – April 2nd

Astronomy Report

Name: Amanda Manget

Crew: 208


Date: 02-APR-2019

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: MDRS-WF

Objects to be Imaged this Evening (if weather permits):
Submitted March 27th: Pinwheel (active 7/10 images taken)
Submitted April 1st: Crab Nebula, Dumbbell Nebula

Objects Already Imaged: Pinwheel with MDRS-14, Whirlpool Galaxy with MDRS-WF



Problems Encountered: Could not identify whirlpool galaxy in the images, even after processing.

MUSK OBSERVATORY



Solar Features Observed: N/A



Problems Encountered: N/A

Cheers,

Julielynn Wong, MD, MPH, FACPM
Chief Medical Maker

Journalist Report – April 2nd

Crew 208 Journalist Report

Date: 2-April-2019

Author: Erika Rydberg

Where is Windy City? Mars!

We began today with some new and exciting food supplies and consumed some new food including: oatmeal, chocolate with almonds, and cheese crackers. We’ve been loving the variety of freeze dried supplies that we have, and love to eat not just fruit, but our broccoli florets as they are — freeze dried. Now onto our mission…

Today with the wind picking up, we cancelled one of our EAVs to continue to test our drone controllers. Amanda, Dean and David were able to go on an expedition and complete a 2nd day of testing as planned prior to the gusting wind. Around the Hab we perfected our 3D drone maps after some continued troubleshooting – it is my belief this will be our best map yet.

I, Crew Journalist, took a ride on the bicycle for one hour. We’re using our bicycle to test power generation to power our 3D printers. I envisioned being outside on a beautiful day with the sun shining down on Mars, while cycling away inside and generating sufficient power for our battery.

We’re finishing the day with some more testing of our thoracentesis trainer and some dinner of jambalaya another item from our new supplies. More to come tomorrow, on what will hopefully be a less windy day.

Cheers,

Julielynn Wong, MD, MPH, FACPM
Chief Medical Maker

GreenHab Report – April 2nd

Crew 208 GreenHab Report – 02-April-2019

GreenHab Officer: Erika Rydberg

Environmental control:

Ambient, door closed.

Shade cloth (80%) on

Average temperatures:

Average: 24.1 C

Low: 27.7
High: 22.6 C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank— 208 gallons

Time(s) of watering for crops: 8:30 and 17:30

Changes to crops: Tomato plant is flowering. Basil is also flowering, I removed the flower buds to keep the basil growing.

Narrative: Cooler day in the Greenhab, the sage was delicious in a goulash-like rice and beans lunch.

Harvest: 5 g sage

Support/supplies needed: None

Cheers,

Julielynn Wong, MD, MPH, FACPM
Chief Medical Maker